A wave has an amplitude of 13.44 cm and a frequency of 284 Hz. Determine the magnitude of the transverse velocity, giving your answer in m/s
Answer #1


given,

wave amplitude A=13.44 cm

frequency f=284 Hz

now,

transverse velocity v=A*w=A*(2pi*f)

v=(13.44*10^-2)*(2pi*284)

v=239.705 m/sec